OpenStudy (anonymous):

z/1 -1/z =0 is the answer (z,-z)

OpenStudy (anonymous):

\[\frac{z}{1}-\frac{1}{z}=0\] denominator is (1*z) \[z*z= z^{2}\] \[(z^{2}-1)=o\] factor: (x+1)(x-1) =0 z = 1 z = -1 If I deserve medal then give me one.

OpenStudy (anonymous):

how do I do that, I am new to this

OpenStudy (anonymous):

next to my name click give medal

OpenStudy (anonymous):

di u se/

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

I think I did it

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

yes, any more question?

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

post it I will try the best

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

find the LCM y^3 +4y^2 + 4y, y^2-4y

OpenStudy (anonymous):

2 question?

OpenStudy (anonymous):

no, thats 1

OpenStudy (anonymous):

should i post it on the otherside so u can get credit for this as well

OpenStudy (anonymous):

\[y^{3}+4y ^{2}+4y \] and another on is : \[y^{2}-4y\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

look like u have 2 not 1 ?

OpenStudy (anonymous):

the question gives 5 numbers and ask me to find the LCM the way it wants me to answer is number (number)^exponent (y-number)

OpenStudy (anonymous):

LCM = 4y

OpenStudy (anonymous):

y^3 +4y^2 + 4y= 4y(y^2 +y+1) y^2-4y = 4y(y-1) than LCM is ; 4y

OpenStudy (anonymous):

wailt

OpenStudy (anonymous):

I'm not sure u may re post question to see some one beter

OpenStudy (anonymous):

I can factor nut not sure LCM

OpenStudy (anonymous):

no prroblem thanks for trying, I do have something that you may be able to help with

OpenStudy (anonymous):

y^3 +4y^2 + 4y = y(y^2+4y+4) = y(y+2)(y+2) =y (y+2)^2 y=0 , y=-2

OpenStudy (anonymous):

y^2-4y = y(y-4) y = 0 , y =4

OpenStudy (anonymous):

here goes another one x^2-3x+2/x^2+5x-6

OpenStudy (anonymous):

it states that i should simplremoving factors of one

OpenStudy (anonymous):

I can do this ,give a time I'm slow

OpenStudy (anonymous):

ok, lol no problem, and thank u

OpenStudy (anonymous):

\[(x^{2}-3x+2)\div(x ^{2}+5x -6)\] I want made suse I'm right?

OpenStudy (anonymous):

no not division its factoring

OpenStudy (anonymous):

\[\frac{(x-2)(x-1)}{(x+3)(x-2)}\]

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

x- 2 cancel\[\frac{(x-1)}{(x+3}\]

OpenStudy (anonymous):

that right?

OpenStudy (anonymous):

i believe so, im horrible at math

OpenStudy (anonymous):

I sure it correct

OpenStudy (anonymous):

ok, thanx

OpenStudy (anonymous):

u r welcom